express the decimal number 4693 as a binary number
Found 2 solutions by stanbon, checkley77:
Answer by stanbon(75887)

express the decimal number 4693 as a binary number
-----------
Where is the decimal point?
--------
Cheers,
Stan H.

Answer by checkley77(12844)

4,693=1001001010101
Proof:
4096*1=4096
2048*0=0
1024*0=0
512*1=512
256*0=0
128*0=0
64*1=64
32*0=0
16*1=16
8*0=0
4*1=4
2*0=0
1*1=1
----------
4,693